Part 3..PHY2...16
A conducting ring is in a uniform magnetic field B
Answer #1

given that,

radius,r=50 cm=0.05 m

resistance,R=0.10 ohm

magnetic field, B0=0.042 T

time, t=0.20 s

a)

e=-N\Delta \phi/\Deltat

=-BA/t

=Bpir2/t

=(0.042 T*3.14*0.05 m*0.05 m)/(0.20 s)

=0.16485 V

b)

e=IR

I=e/R

=(0.16485 V)/(0.10 ohm)

=1.6485 A